Find the molar concentration of 5 mg/mL solution of Hemoglobin and Myoglobin
Answer #1

Solution :-

5 mg/mL

5 mg * 1 g / 1000 mg= 0.005 g

Molar mass of Hemoglobin = 65000 g/mol

Molar mass of myoglobin = 17600 g/mol

Using the mass and molar mass we can find the moles

Moles= mass/ molar mass

Moles of Hemoglobin = 0.005 g / 65000 g per mol

                                       = 7.7*10^-8 mol

Moles of myoglobin = 0.005 g / 17600 g per mol

                                    = 2.8*10^-7 mol

Molarity = moles / volume in liter

1 mL*1 L / 1000 mL= 0.001 L

Molarity of hemoglobin = 7.7*10^-8 mol / 0.001 L

                                            = 7.7*10^-5 M

Molarity of myoglobin = 2.8*10^-7 mol / 0.001 L

                                         = 2.8*10^-4 M